# CERN-TH-2024-065 ET-0343A-24

### v1

@@ V.2.2 Heavy QCD axions
As usual, annihilation of the string-wall network is determined by the aforementioned explicit breakings of the PQ symmetry, which can be captured by an additional misaligned potential $V_{b}\simeq-\mu_{b}^{4}\cos\left(\frac{N_{\text{b}}}{N_{\text{DW}}}\frac{a}{F_{a}}-\delta\right),$ where $N_{b}$ should be $1$ or co-prime with $N_{\text{DW}}$ in order to induce annihilation. For illustration we take the representative values $N_{b}=1,\kappa_{\text{H}}=1$ and $N_{\text{DW}}=6,\delta=0.3$ in the following (see Ferreira et al. (2023b) for other choices). The annihilation temperature is then
@@
V.2.3 FOPT at ground-based detectors + DW signal at PTA
@@ VI Conclusion
The aim of this review was to quantify the ability of 3G ground based detectors, working together with LISA and SKA, to probe cosmological SGWBs sourced by PTs in the early universe. PT and symmetry breaking are ubiquitous in quantum field theories and particle physics beyond the SM, and as we have discussed they can lead to copious GW production across different frequency bands. Among the most interesting possible sources of GWs are (i) first order PTs as reviewed in section III, and (ii) topological defects of different kinds, as reviewed in section IV. Figure 2 shows examples of characteristic spectra from these sources, together with the PLISCs of the detectors across the frequency band, accounting for foregrounds.
@@
In section III, we have focused on GW production by a first order PT. We have first reviewed the PT dynamics and defined the relevant stages of its progression, and the associated parameters such as the PT duration and the size of the bubbles at percolation. We have then reviewed the phenomena associated with the PT capable of sourcing a SGWB, namely: bubbles and/or relativistic fluid shells collisions for strongly first order PTs; bulk fluid motion generated at percolation in the form of sound waves and (M)HD turbulence for PTs of intermediate strength; and sound waves alone for weak PTs. We have shown how the scaling of the SGWB signals with the source parameters, provided by detailed analyses of the aforementioned sourcing mechanisms, can be interpreted based on simple models for the source UETC, namely, constant-in-time and stationary. We have then presented placeholders for the SGWB spectra from the three generation mechanisms, highlighting the assumptions behind them. Evaluating the SGWB signal from a first order PT is a complicated endeavour, requiring particle physics and statistical physics as well as non-equilibrium phenomena, fluid dynamics and cosmology, and further progress is needed to clarify and overcome all the caveats that enter in this process. Nevertheless, we have concluded by showing the reach of future 3G detectors (as well as LISA and PTA with the SKA) to the compelling case of SGWBs generated by first order PTs, based on the state of the art of our knowledge. The result is presented in Fig. 3 as a scan on the SNR of the SGWB signal in CE, ET, LISA and SKA, as a function of the parameter space of first order PTs (characterised by the transition temperature $T_{*}$ and its duration $\beta/H_{*}$, for different fixed strengths $\alpha=0.05,1,\gg 1$), including the astrophysical foregrounds. The figure clearly shows the potential of these detectors to probe first order PTs in full complementarity with respect to the PT energy scale (represented by its temperature). It also shows that, in a large region of parameter space, the signal will have a higher SNR than that of the expected astrophysical foregrounds. For strong first order PTs ($\alpha\gg 1$), some regions of parameter space are already excluded by LVK O3 data, and others could account for the observed PTA signal.
@@
In section IV we have focused on topological defects which may be formed in PTs provided a symmetry is broken. Their existence depends on the properties of the vacuum manifold ${\cal{M}}$, and the argument for defect formation is based solely on causality: in other words, they may form whether the PT is first order or not. For local $U(1)$ strings, whose tension is given by $\eta^{2}$, where $\eta$ the energy scale of the PT, we reviewed in section IV.1 the shape of the spectrum which covers orders of magnitude in frequency, and whose amplitude is proportional to $(G\mu)^{2}$. As discussed, understanding the detailed properties of the spectral shape still require further theoretical understanding. In that section we have shown that, with a precise estimation of the astrophysical foreground, CE should be able to probe strings with tensions $G\mu\gtrsim 10^{-16}$ (Model A) and $G\mu\gtrsim 10^{-18.7}$ (Model B). For ET, one has respectively $G\mu\gtrsim 10^{-14.5}$ and $G\mu\gtrsim 10^{-17.4}$. Both ET and CE will be able to search for individual bursts emitted by cosmic string cusps and kinks, but if none are detected, then constraints on the tension of cosmic loops is less competitive than those from the SGWB.
@@
In section V we then focused on the PQ mechanism as a well-motivated beyond the SM scenario that can potentially be probed by ET and CE (as well as LISA). GWs can arise in this mechanism from the PQ phase transition, if it is of first order, and/or from its topological byproduct, that is the axionic string-domain wall network. First, we reviewed UV realizations where the PQ sector is approximately scale-invariant, within weakly or strongly coupled frameworks. The PT is then expected to be of first order, and typically exhibits a significant amount of supercooling. The corresponding GW signal is thus mostly sourced by bubble and/or relativistic shells collisions (although bulk fluid motion can be efficient in certain regions of the parameter space) and is particularly loud at ET/CE, since the scale of the PT $\sim F_{a}$ is constrained to be around $10^{8}-10^{10}~{}\text{GeV}$ from cosmological and astrophysical bounds on the QCD axion. We presented a forecast of the regions of parameter space in strongly-coupled PQ models which will be probed by 3G detectors in Fig. 7, whereas we presented spectra for benchmark points in the case of weakly coupled models. Our analysis improves significantly over previous work Von Harling et al. (2020), in that we have taken into account more stringent conditions on the completion of PTs in vacuum dominated eras, beside the calculation of the SNR of the signals. In both cases, we have shown that significant regions of parameter space will be probed by 3G detectors. This provides the exciting opportunity to explore the UV implementation of the PQ mechanism, for values of $F_{a}$ where the QCD axion can be the Dark Matter.
@@
Secondly and independently of the order of the PQ phase transition, we have investigated GWs from axionic domain walls. While the signal is too small to be observable in the standard QCD axion scenario, we showed that DWs can source a large GW signal in PQ models where the axion is much heavier than usual, as can occur in UV models that include additional contributions to the QCD potential. While UV realizations of these models require additional ingredients, the scenario can potentially alleviate the axion quality problem, while also leading to heavy DWs and large GW backgrounds. The relevant GW abundance and frequencies depend on the axion mass and decay constant, and additionally on the scale at which the axionic discrete symmetry is broken to allow a viable cosmology without DWs at late times. Our results are presented in Fig. 8 and show that the GW signal is louder than astrophysical foregrounds in a large region of parameter space of these models. Very interestingly, the region that can be probed via GWs in Fig. 8 also leads to observable electric dipole moments for the neutron and the proton, thereby providing a complementary channel to test heavy axion models.
Here we have focused on PTs as a well understood and physically well motivated phenomenon that sources GWs. However, perhaps the most exciting perspective is that, by probing early Universe temperatures as high as $\gtrsim 10^{8}~{}\text{GeV}$, 3G detectors may reveal imprints of totally new phenomena which occurred in the primordial universe.
@@ Appendix A Overlap Reduction Function
$h(t,\vec{x})$ $=\int_{-\infty}^{+\infty}\mathrm{d}\tau\int\mathrm{d}^{3}yR^{\mu\nu}(\tau,\vec{y})h_{\mu\nu}(t-\tau,\vec{x}-\vec{y})$ (117)
@@ Appendix B Einstein Telescope SNR computation
Contrary to the usual case of uncorrelated detector noises, one cannot simply add up the SNR squared of the individual detector pairs as their associated statistics is clearly correlated from Eq. (130). Instead, one can derive the way to linearly combine these estimators in order to minimize the variance of the resulting statistics while preserving their expectation value (unbiased estimator). This is a well known problem already derived in Allen and Romano (1999); Romano and Cornish (2017). To simplify the notation, we denote with $A$ and $B$ interferometer pairs $(I,J)$ and $(K,L)$ (note again that $A$ and $B$ do not need to be distinct) so that $\hat{\mathcal{S}}_{A}=\hat{\mathcal{S}}_{IJ}$. We now define a statistic that linearly combines the one from individual interferometer pairs $A$
@@
where we write the SNR for a general number of $N$ equivalent colocated interferometers. For ET, $N=3$. Using the usual method from Allen and Romano (1999); Romano and Cornish (2017), one can choose the filter function $\tilde{Q}$ to maximize the SNR. One finds

### v2

@@ V.2.2 Heavy QCD axions
The annihilation of the string-wall network is determined by the aforementioned explicit breakings of the PQ symmetry, which can be captured by an additional misaligned potential $V_{b}\simeq-\mu_{b}^{4}\cos\left(\frac{N_{\text{b}}}{N_{\text{DW}}}\frac{a}{F_{a}}-\delta\right),$ where $N_{b}$ should be $1$ or co-prime with $N_{\text{DW}}$ in order to induce annihilation. For illustration we take the representative values $N_{b}=1,\kappa_{\text{H}}=1$ and $N_{\text{DW}}=6,\delta=0.3$ in the following (see Ferreira et al. (2023b) for other choices). The annihilation temperature is then
@@
V.2.3 First order PT at ground-based detectors + DW signal at PTA
@@ VI Conclusion
The aim of this review was to quantify the ability of 3G ground based detectors, working together with LISA and SKA, to probe cosmological SGWBs sourced by PTs in the early universe. PT and symmetry breaking are ubiquitous in quantum field theories and particle physics beyond the SM and, as we have discussed, they can lead to copious GW production across different frequency bands. Among the most interesting possible sources of GWs are (i) first order PTs as reviewed in section III, and (ii) topological defects of different kinds, as reviewed in section IV. Figure 2 shows examples of characteristic spectra from these sources, together with the PLISCs of the detectors across the frequency band, accounting for foregrounds. The broad and complementary frequency span of 3G ground based detectors, LISA, and PTA, will be a fundamental asset for probing the variety of SGWB signals that can arise from PTs in the early Universe. For example, the coincident detection of a SGWB from the same generation process across the frequency bands would offer enhanced information on the process itself, and would fortify its interpretation as a primordial signal. Furthermore, the wider the frequency coverage, the broader the time span in the early Universe that one can access, see Eq. 2. Fig. 2 clearly shows that the combination of future GW detectors opens a new window of exploration of the early Universe, spanning 11 decades in frequency and therefore in energy. However, Fig. 2 also shows that the presence of foregrounds highly influences the sensitivity of some detectors, the worst case being PTA with the SKA, for which accounting for the SMBHB foreground changes the sensitivity by about 3 orders of magnitude.
@@
In section III, we have focused on GW production by a first order PT. We have first reviewed the PT dynamics and defined the relevant stages of its progression, and the associated parameters such as the PT duration and the size of the bubbles at percolation. We have then reviewed the phenomena associated with the PT capable of sourcing a SGWB, namely: bubbles and/or relativistic fluid shells collisions for strongly first order PTs; bulk fluid motion generated at percolation in the form of sound waves and (M)HD turbulence for PTs of intermediate strength; and sound waves alone for weak PTs. We have shown how the scaling of the SGWB signals with the source parameters, provided by detailed analyses of the aforementioned sourcing mechanisms, can be interpreted based on simple models for the source UETC, namely, constant-in-time and stationary. We have then presented placeholders for the SGWB spectra from the three generation mechanisms, highlighting the assumptions behind them. Evaluating the SGWB signal from a first order PT is a complicated endeavour, requiring particle physics and statistical physics as well as non-equilibrium phenomena, fluid dynamics and cosmology, and further progress is needed to clarify and overcome all the caveats that enter in this process. Nevertheless, we have concluded by showing the reach of future 3G detectors (as well as LISA and PTA with the SKA) to the compelling case of SGWBs generated by first order PTs, based on the state of the art of our knowledge. The result is presented in Fig. 3 as a scan on the SNR of the SGWB signal in CE, ET, LISA and SKA, as a function of the parameter space of first order PTs (characterised by the transition temperature $T_{*}$ and its duration $\beta/H_{*}$, for different fixed strengths $\alpha=0.05,1,\gg 1$), including the astrophysical foregrounds. The figure clearly shows the potential of the considered detectors to probe first order PTs in full complementarity with respect to the PT energy scale (represented by its temperature). Therefore, when all these observatories will be in operation, they will offer the opportunity to probe a wide range of energy scales in the early Universe. 3G detectors allow to explore new, high energy scales for first order PTs with respect to the PTs predicted within the SM, probed rather by LISA (EW scale) and PTA (QCD scale). Particularly interesting is the possibility to probe relatively weak ($\alpha=0.05$) and fast ($\beta/H_{*}\simeq 100-1000$) PTs, predicted by many BSM models. Fig. 3 also shows that, in a large region of parameter space, the signal will have higher SNR than that of the expected astrophysical foregrounds. For strong first order PTs ($\alpha\gg 1$), some regions of parameter space are already excluded by LVK O3 data, and others could account for the observed PTA signal.
@@
In section IV we have focused on topological defects which may be formed in PTs provided a symmetry is broken. Their existence depends on the properties of the vacuum manifold ${\cal{M}}$, and the argument for defect formation is based solely on causality: in other words, they may form whether the PT is first order or not. For local $U(1)$ strings, whose tension is given by $\eta^{2}$, where $\eta$ the energy scale of the PT, we reviewed in section IV.1 the shape of the spectrum which covers orders of magnitude in frequency, and whose amplitude is proportional to $(G\mu)^{2}$. As discussed, understanding the detailed properties of the spectral shape still requires further theoretical understanding. In that section we have shown that, with a precise estimation of the astrophysical foreground, CE should be able to probe strings with tensions $G\mu\gtrsim 10^{-16}$ (Model A) and $G\mu\gtrsim 10^{-18.7}$ (Model B). For ET, one has respectively $G\mu\gtrsim 10^{-14.5}$ and $G\mu\gtrsim 10^{-17.4}$. Both ET and CE will be able to search for individual bursts emitted by cosmic string cusps and kinks, but if none are detected, then constraints on the tension of cosmic loops is less competitive than those from the SGWB.
@@
In section V we then focused on the PQ mechanism as a well-motivated beyond the SM scenario that can potentially be probed by ET and CE (see also Yang and Huang (2024)), as well as LISA. GWs can arise in this mechanism from the PQ phase transition, if it is of first order, and/or from its topological byproduct, that is the axionic string-domain wall network. First, we reviewed UV realizations where the PQ sector is approximately scale-invariant, within weakly or strongly coupled frameworks. The PT is then expected to be of first order, and typically exhibits a significant amount of supercooling. The corresponding GW signal is thus mostly sourced by bubble and/or relativistic shells collisions (although bulk fluid motion can be efficient in certain regions of the parameter space) and is particularly loud at ET/CE, since the scale of the PT $\sim F_{a}$ is constrained to be around $10^{8}-10^{10}~{}\text{GeV}$ from cosmological and astrophysical bounds on the QCD axion. We presented a forecast of the regions of parameter space in strongly-coupled PQ models which will be probed by 3G detectors in Fig. 7, whereas we presented spectra for benchmark points in the case of weakly coupled models. Our analysis improves significantly over previous work Von Harling et al. (2020), in that we have taken into account more stringent conditions on the completion of PTs in vacuum dominated eras, beside the calculation of the SNR of the signals. In both cases, we have shown that significant regions of parameter space will be probed by 3G detectors. In particular, the GW signal is loudest at 3G detectors when the axion decay constant lies towards the lower end of the traditional QCD axion window, i.e. $F_{a}\sim 10^{8}-10^{9}~{}\text{GeV}$, although an observable signal can be obtained also for larger decay constants. This corresponds to QCD axion masses in the range $m_{a}\sim\text{meV}-0.1~{}\text{eV}$. For significantly lighter masses the signal would be peaked at frequencies larger than the detectors sensitivity regions. Concerning dark matter, the detection of a SGWB from the early Universe at 3G detectors, if interpreted as coming from the PQ mechanism, would suggest that the axion decay constant is compatible with the QCD axion being the Dark Matter. In this case, the Dark Matter would be produced either by cosmic strings and the misalignment mechanism (for $F_{a}\lesssim 10^{9}-10^{10}~{}\text{GeV}$) or by domain walls (for smaller values).
@@
Secondly and independently of the order of the PQ phase transition, we have investigated GWs from axionic domain walls. While the signal is too small to be observable in the standard QCD axion scenario, we showed that DWs can source a large GW signal in heavy QCD axion models, where the axion mass is an additional independent parameter. While UV realizations of these models require additional ingredients, the scenario can potentially alleviate the axion quality problem, while also leading to heavy DWs and large GW backgrounds. The relevant GW abundance and frequencies depend on the axion mass and decay constant, and additionally on the scale at which the axionic discrete symmetry is broken to allow a viable cosmology without DWs at late times. Our results are presented in Fig. 8 and show that the GW signal is louder than astrophysical foregrounds in a large region of parameter space of these models. In particular, the GW signal from domain walls is detectable at 3G detectors when $10^{10}~{}\text{GeV}\lesssim F_{a}\lesssim 10^{11}~{}\text{GeV}$, and the axion is very heavy, $10^{6}~{}\text{GeV}\lesssim m_{a}\lesssim 10^{11}~{}\text{GeV}$. The signal from domain walls is louder for larger values of the mass and decay constant. Very interestingly, the region that can be probed via GWs in Fig. 8 also leads to observable electric dipole moments for the neutron and the proton, thereby providing a complementary channel to test heavy axion models. Concerning axion dark matter, the important difference from the first order PT scenario is that the detection of a GW signal from DWs at 3G detectors, together with the observation of electric dipole moments, would give support to scenarios where the axion is unstable and thus cannot be the dark matter.
The possibility of observing GW backgrounds is informative even in the case of null detection. For the case of the QCD axion, this translates into constraining realizations of the PQ mechanism which predict a first order PT (the simplest being scenarios with approximate scale invariance reviewed above) and heavy axion models with long-lived domain walls, as can arise when $N_{\text{dw}}>1$. Existing observations from LVK have already started constraining those scenarios Badger et al. (2023); Jiang and Huang (2022).
Here we have focused on PTs as a well understood and physically well motivated phenomenon that sources GWs. However, perhaps the most exciting perspective is that, by probing early Universe temperatures as high as $\gtrsim 10^{8}~{}\text{GeV}$, 3G detectors may reveal imprints of totally new phenomena which occurred in the primordial universe. As summarized over the review, reaching this goal will require improving the characterization and understanding of the instruments and the astrophysical foregrounds. In parallel, the characterization of the primordial sources, which often require heavy numerical computations, still requires significant improvement. This includes the understanding of the associated signals, in the form of SGWBs in various observatories, but also possibly as dark matter, dark radiation, primordial black holes or even electric dipole moments as in the case of axions. These additional signals will of course play a crucial role in model discrimination.
@@ Appendix A Overlap Reduction Function
$h(t,\vec{x})$ $=\int_{-\infty}^{+\infty}\mathrm{d}{t^{\prime}}\int\mathrm{d}^{3}yR^{\mu\nu}({t^{\prime}},\vec{y})h_{\mu\nu}(t-{t^{\prime}},\vec{x}-\vec{y})$ (117)
@@ Appendix B Einstein Telescope SNR computation
Contrary to the case of uncorrelated detector noises, one cannot simply add up the SNR squared of the individual detector pairs as their associated statistics is clearly correlated from Eq. (130). Instead, one can derive the way to linearly combine these estimators in order to minimize the variance of the resulting statistics while preserving their expectation value (unbiased estimator). This is a well known problem already derived in Allen and Romano (1999); Romano and Cornish (2017). To simplify the notation, we denote with $A$ and $B$ interferometer pairs $(I,J)$ and $(K,L)$ (note again that $A$ and $B$ do not need to be distinct) so that $\hat{\mathcal{S}}_{A}=\hat{\mathcal{S}}_{IJ}$. We now define a statistic that linearly combines the one from individual interferometer pairs $A$
@@
where we write the SNR for a general number of $N$ equivalent colocated interferometers. For ET, $N=3$. Using the standard method from Allen and Romano (1999); Romano and Cornish (2017), one can choose the filter function $\tilde{Q}$ to maximize the SNR. One finds